OpenStudy (anonymous):

x/5-g=a for x someone help!!!!!

OpenStudy (anonymous):

\[\frac{x}{5}-g=a\]?

OpenStudy (anonymous):

or \[\frac{x}{5-g}=a\]

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

add \(g\) then multiply by \(5\)

OpenStudy (anonymous):

when i say "add" and "multiply" i just mean " write it" \[\frac{x}{5}-g=a\\ \frac{x}{5}=a+g\\ x=5(a+g)\]

OpenStudy (anonymous):

thanks !!!

OpenStudy (anonymous):

yw